ANLA List

AxNey03a [1] O. Axelsson and M. Neytcheva, Preconditioning methods for linear systems arising in constrained optimization problems, Numer. Linear Algebra Appl., 10 (2003), pp. 3–31.

AxNey06 [2] , Eigenvalue estimates for preconditioned saddle point matrices, Numer. Linear Algebra Appl., 13 (2006), pp. 339–360.

BaNey05b [3] E. Bängtsson and M. Neytcheva, Numerical simulations of glacial rebound using preconditioned iterative solution methods, Appl. Math., 50 (2005), pp. 183–201.

BlMarNey05 [4] R. Blaheta, S. Margenov, and M. Neytcheva, Robust optimal multilevel preconditioners for non-conforming finite element systems, Numer. Linear Algebra Appl., 12 (2005), pp. 495–514.

BrHoSu07 [5] H. Brandén, S. Holmgren, and P. Sundqvist, Discrete fundamental solution preconditioning for hyperbolic systems of PDE, J. Sci. Comput., 30 (2007).

BrSu05 [6] H. Brandén and P. Sundqvist, Preconditioners based on fundamental solutions, BIT, 45 (2005), pp. 481–494.

GusHe99 [7] B. Gustafsson and L. Hemmingsson-Frändén, A fast domain decomposition high order Poisson solver, J. Sci. Comput., 14 (1999), pp. 223–243.

Hemm96a [8] L. Hemmingsson, Toeplitz preconditioners with block structure for first-order PDEs, Numer. Linear Algebra Appl., 3 (1996), pp. 21–44.

Hemm98 [9] , A semi-circulant preconditioner for the convection-diffusion equation, Numer. Math., 81 (1998), pp. 211–248.

HeOtto96 [10] L. Hemmingsson and K. Otto, Analysis of semi-Toeplitz preconditioners for first-order PDEs, SIAM J. Sci. Comput., 17 (1996), pp. 47– 64.

HoBrSt00 [11] S. Holmgren, H. Brandén, and E. Sterner, Convergence acceleration for the linearized Navier–Stokes equations using semicirculant approximations, SIAM J. Sci. Comput., 21 (2000), pp. 1524–1550.

HoOtto92 [12] S. Holmgren and K. Otto, Iterative solution methods and preconditioners for block-tridiagonal systems of equations, SIAM J. Matrix Anal. Appl., 13 (1992), pp. 863–886.

HoOtto94 [13] , Semicirculant preconditioners for first-order partial differential equations, SIAM J. Sci. Comput., 15 (1994), pp. 385–407.

HoOtto96 [14] , Semicirculant solvers and boundary corrections for first-order partial differential equations, SIAM J. Sci. Comput., 17 (1996), pp. 613– 630.

HoOtto98a [15] , A framework for polynomial preconditioners based on fast transforms I: Theory, BIT, 38 (1998), pp. 544–559.

HoOtto98b [16] , A framework for polynomial preconditioners based on fast transforms II: PDE applications, BIT, 38 (1998), pp. 721–736.

Larsson99 [17] E. Larsson, A domain decomposition method for the Helmholtz equation in a multilayer domain, SIAM J. Sci. Comput., 20 (1999), pp. 1713–1731.

LaHolm03 [18] E. Larsson and S. Holmgren, Parallel solution of the Helmholtz equation in a multilayer domain, BIT, 43 (2003), pp. 387–411.

LSRH02 [19] P. Lötstedt, S. Söderberg, A. Ramage, and L. Hemmingsson-Frändén, Implicit solution of hyperbolic equations with space-time adaptivity, BIT, 42 (2002), pp. 134–158.

Otto96 [20] K. Otto, Analysis of preconditioners for hyperbolic partial differential equations, SIAM J. Numer. Anal., 33 (1996), pp. 2131–2165.

OttoLa99 [21] K. Otto and E. Larsson, Iterative solution of the Helmholtz equation by a second-order method, SIAM J. Matrix Anal. Appl., 21 (1999), pp. 209–229.

AhlMun05 [22] K. Åhlander and H. Munthe-Kaas, Applications of the generalized Fourier transform in numerical linear algebra, BIT, 45 (2005), pp. 819–